3. Make sure you answer all three parts.
A square floor tile has an area of (x2 + 8x +16) in2. The side length of the tile is of the form cx + d, where c and d are whole numbers.
a. Find the expression for the side length of the tile.
b. Find the expression that represents the perimeter of the tile.
c. Find the perimeter when x=8 inches.

Answer by ewatrrr(24785)

(x^2 + 8x +16) in^2 = (x+4)^2 in^2
........
side length: s = (x+4)
P = 4s
P = 4(x+4)in
......
x = 8
P = 4(8+4)in = 48in